OpenStudy (anonymous):

Use the Special Right Triangle to evaluate sin 45°, cos 45° and tan 45°. This question has already been posted but it wasn't explained how they got the answer.

OpenStudy (anonymous):

You will have a ninety degree angel and two 45. The rule for 45 special triangles is

OpenStudy (anonymous):

Sorry, @nelsonjedi I still don't understand very well :/. Am I on the right track with this? Sine=opposite/hypotenuse sine=x/x(sqrt{2}) Cosine=adjacent/hypotenuse cosine=x/x(sqrt{2}) Tangent=opposite/adjacent tangent=x/x

OpenStudy (anonymous):

Winner winner chicken dinner,,,great job now you need to simplify....

OpenStudy (anonymous):

Alrighty... Sine=x^2/sqrt(2) Cosine=x^2/sqrt(2) Tangent=1 Is that simplified? Never been good at simplification.

OpenStudy (anonymous):

The x values cael each other out...the we multiply the top and bottom by sq root of 2 get get the irrational out \[\frac{ x }{ x \sqrt{2}} = \frac{ 1 }{ \sqrt{2} } = \frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

Thank you so much! You were a great help @nelsonjedi